OpenStudy (anonymous):

Please help! Solve for x: 3/4x+5/8=4x

OpenStudy (anonymous):

x = 0.19

OpenStudy (anonymous):

write the equation with a clearer notation please. is it \[\frac{ 3 }{ 4 }x,\frac{ 3 }{ 4x }\]

OpenStudy (anonymous):

@nerdtastic365 Does the equation look like this?\[\frac{ 3 }{ 4 }x + \frac{ 5 }{ 8 } = 4x\]

OpenStudy (anonymous):

If it does then x = 0.19

OpenStudy (anonymous):

Yes, sorry for not being more clear @greenlegodude57 and @VeritasVosLiberabit

OpenStudy (anonymous):

But, I would just like to know the steps to solving the problem, please, not the answer